OpenStudy (dragon123mama):

Suppose a Parabola has an axis of symmetry at x= -5 , a mazimum hight of 9 , and passes through the point (-7 , 1) what is the equation in vertex form 1. y= -4(x-5)^2 +9 2. y= -7(x+9)^s -5 3.y= -0.06 (x-5)^2 +9 4. y= -2(x+5)^2 +9

OpenStudy (welshfella):

the vertex is at (-5, 9) and is a maximum the vertex form is a(x - b)^2 - c where (b,c) is the vertex and a is some constant

OpenStudy (dragon123mama):

so which one is it then? 1234? i guessed 4 but i am not sure.

OpenStudy (welshfella):

we have b = -5 and c = 9 you can find a by plugging in the point (-7,1) and solving for a.

OpenStudy (welshfella):

y = a( x + 5)^2 - 9 now plug in x = -7 and y = 1 and solve for a.

OpenStudy (dragon123mama):

so it is 1 or 3...?

OpenStudy (welshfella):

find a and you'll know

OpenStudy (dragon123mama):

4?

OpenStudy (dragon123mama):

choice 4?

OpenStudy (dragon123mama):

none of the choices have -9 at the end....?

OpenStudy (welshfella):

oh apologies its y = a(x + 5)^2 + 9 (not -9)

OpenStudy (dragon123mama):

u said it was -5 as well....? didn't you earlier..?

OpenStudy (dragon123mama):

so would it be 1...? option 1.

OpenStudy (welshfella):

yes that is the value of x at the vertex but standard form is a(x - b)^2 + c and b = -5

OpenStudy (welshfella):

y = a(x + 5)^2 + 9 put y = 1 and x = -7 then solve for a

OpenStudy (dragon123mama):

so it is option 4. okay.

OpenStudy (welshfella):

show your work please

OpenStudy (dragon123mama):

i did it in my head. i am just doing a review for my study guide. nothing important. my mom made these equations for my to study before my test. i just need people to check my answeres. so it was 4. i was right.